# Adaptive Online Bayesian Estimation of Frequency Distributions with Local Differential Privacy

### v1

@@ A.4.2 Convergence of the posterior distribution
for random variables $(\Theta\in\Delta,\{S_{t}\subset\{1,\ldots,K\},Y_{t}\in\{1,\ldots,K\}\}_{t\geq 1})$, where $\mathrm{d}\theta$ is the Lebesgue measure for $\vartheta$ restricted to $\Delta$ and $|\Delta|:=\int_{\Delta}\mathrm{d}\theta$. We will show that the parameter $\mu$ in (49) can be chosen such that the collection of random variables
@@
$E^{\otimes}[e^{nm_{0}\mathcal{E}^{\mu}_{n}(\Theta,\theta^{\ast})}\mathbb{I}(f_{n}>K)]$ $\leq\sup_{\theta\in A_{n}}e^{nm_{0}\mathcal{E}^{\mu}_{n}(\theta,\theta^{\ast})}P^{\otimes}(f_{n}>K)$
$=\sup_{\theta\in A_{n}}e^{nm_{0}\mathcal{E}^{\mu}_{n}(\theta,\theta^{\ast})}\int_{\Delta}P_{\theta^{\ast}}\left(\mathcal{E}^{\mu}_{n}(\theta,\theta^{\ast})>\frac{\ln K}{nm_{0}}\right)\frac{\mathrm{d}\theta}{|\Delta|}$
@@
The parameter $\mu$ can be arranged such that (51) converges to 00. For such $\mu$, we have that for any $\varepsilon$ there exists a $N_{\varepsilon}>0$ such that for all $n>N_{\varepsilon}$, $E^{\otimes}[|f_{n}|\cdot\mathbb{I}(f_{n}>K)]<\varepsilon$ for any $K>0$. Finally, choose $K_{\varepsilon}=\sup_{\theta\in A_{N_{\varepsilon}}}e^{N_{\varepsilon}m_{0}\mu c_{u}}$ so that $E^{\otimes}[|f_{n}|\cdot\mathbb{I}(f_{n}>K_{\varepsilon})]<\varepsilon$ for any $n\geq 1$. Hence, $\mathcal{C}$ is uniformly integrable for a suitable choice of $\mu$.
Next, we show that each $f_{n}$ in $\mathcal{C}$ converges in probability to $1$. The convergence is implied by the fact that for every $\theta\in A_{n}$ and $\varepsilon>0$, we have
@@
that is, the sequence $\int_{A_{n}}\max\{1,e^{nm_{0}\mathcal{E}^{\mu}_{n}(\theta,\theta^{\ast})}\}\mathrm{d}\theta$ converges to $1$ in distribution with respect to $P_{\theta^{\ast}}$. Since convergence in distribution to a constant implies convergence in probability, we have
@@
converges to 00 in probability, which in turn proves the convergence of $\frac{\Pi(A_{n}|Y_{1:n},S_{1:n})}{\Pi(B_{n}|Y_{1:n},S_{1:n})}$ in probability to 00. To do that, we need to prove that for each $\varepsilon>0$ and $\delta>0$ there exists a $N>0$ such that for all $n>N$ we have $P_{\theta^{\ast}}(Z_{n}\geq 2\varepsilon)<2\delta$. Fix $\varepsilon>0$ and $\delta>0$.
@@
  • Next, let $\beta:=C_{1}\max_{S\subset\{1,\ldots,K\}}\textup{Tr}(F(\theta^{\ast};S,\epsilon))/\delta$. Using Markov’s inequality for $\|\nabla_{\vartheta}\Phi_{n}(\theta^{\ast})\|^{2}$ with Lemma 15, we have

### v2

@@ A.4.2 Convergence of the posterior distribution
for random variables $(\Theta\in\Delta,\{S_{t}\subset\{1,\ldots,K\},Y_{t}\in[K]\}_{t\geq 1})$, where $\mathrm{d}\theta$ is the Lebesgue measure for $\vartheta$ restricted to $\Delta$ and $|\Delta|:=\int_{\Delta}\mathrm{d}\theta$. We will show that the parameter $\mu$ in (50) can be chosen such that the collection of random variables
@@
$E^{\otimes}[|f_{n}|\cdot\mathbb{I}(f_{n}>K)]$ $=E^{\otimes}[e^{nm_{0}\mathcal{E}^{\mu}_{n}(\Theta,\theta^{\ast})}\mathbb{I}(f_{n}>K)]$
$\leq\sup_{\theta\in\Delta}e^{nm_{0}\mathcal{E}^{\mu}_{n}(\theta,\theta^{\ast})}P^{\otimes}(f_{n}>K)$
$=\sup_{\theta\in\Delta}e^{nm_{0}\mathcal{E}^{\mu}_{n}(\theta,\theta^{\ast})}\int_{\Delta}P_{\theta^{\ast}}\left(\mathcal{E}^{\mu}_{n}(\theta,\theta^{\ast})>\frac{\ln K}{nm_{0}}\right)\frac{\mathrm{d}\theta}{|\Delta|}$
@@
The parameter $\mu$ can be arranged such that (52) converges to 00. For such $\mu$, we have that for any $\varepsilon$ there exists a $N_{\varepsilon}>0$ such that for all $n>N_{\varepsilon}$, $E^{\otimes}[|f_{n}|\cdot\mathbb{I}(f_{n}>K)]<\varepsilon$ for any $K>0$. Finally, choose $K_{\varepsilon}=e^{N_{\varepsilon}m_{0}\mu c_{u}}$ so that $E^{\otimes}[|f_{n}|\cdot\mathbb{I}(f_{n}>K_{\varepsilon})]<\varepsilon$ for any $n\geq 1$. Hence, $\mathcal{C}$ is uniformly integrable for a suitable choice of $\mu$.
Next, we show that each $f_{n}$ in $\mathcal{C}$ converges in probability to $1$. The convergence is implied by the fact that for every $\theta\in\Delta$ and $\varepsilon>0$, we have
@@
that is, the sequence $\int_{\Delta}\max\{1,e^{nm_{0}\mathcal{E}^{\mu}_{n}(\theta,\theta^{\ast})}\}\mathrm{d}\theta$ converges to $1$ in distribution with respect to $P_{\theta^{\ast}}$. Since convergence in distribution to a constant implies convergence in probability, we have
@@
converges to 00 in probability, which in turn proves the convergence of $\frac{\Pi(A_{n}|Y_{1:n},S_{1:n})}{\Pi(B_{n}|Y_{1:n},S_{1:n})}$ in probability to 00. To do that, we need to prove that for each $\varepsilon>0$ and $\delta>0$, there exists a $N>0$ such that for all $n>N$ we have $P_{\theta^{\ast}}(Z_{n}\geq 2\varepsilon)<2\delta$. Fix $\varepsilon>0$ and $\delta>0$.
@@
  • Next, let $\beta:=C_{1}\max_{S\subset[K]}\textup{Tr}(F(\theta^{\ast};S,\epsilon))/\delta$. Using Markov’s inequality for $\|\nabla_{\vartheta}\Phi_{n}(\theta^{\ast})\|^{2}$ with Lemma 15, we have